A clerk whose salary was ​$305 per week was given a raise of ​$61 per week. What percentage increase did the clerk​ receive
Answer by Theo(13342)

61 / 305 = .2 * 100 = 20%.

the increase was 20% of his salary.

20% of 305 = .2 * 305 = 61.